### v1

# Exploiting the Final Component of Generator Architectures for AI-Generated Image Detection

Abstract
With the rapid proliferation of powerful image generators, accurate detection of AI-generated images has become essential for maintaining a trustworthy online environment. However, existing deepfake detectors often generalize poorly to images produced by unseen generators. Notably, despite being trained under vastly different paradigms, such as diffusion or autoregressive modeling, many modern image generators share common final architectural components that serve as the last stage for converting intermediate representations into images. Motivated by this insight, we propose to “contaminate” real images using the generator’s final component and train a detector to distinguish them from the original real images. We further introduce a taxonomy based on generators’ final components and categorize 21 widely used generators accordingly, enabling a comprehensive investigation of our method’s generalization capability. Using only 100 samples from each of three representative categories, our detector—fine-tuned on the DINOv3 backbone—achieves an average accuracy of 98.83% across 22 testing sets from unseen generators.
1 Introduction
“Last in Line, yet First to Reveal.”
— Inspired by Sherlock Holmes
When examining objects like a careful detective, the last touch often reveals the most, having overwritten the layers of history beneath it. Similarly, given an image and the task of determining whether it is real or AI-generated, it is crucial to examine the distinctive traces left by the generator. Contemporary image generators are no longer monolithic; rather, they comprise multiple modular components operating in stages. This raises a natural and important question: does the last component, which serves as the final step in producing the image, leave identifiable evidence that can be exploited for detection? Our intuition is that, regardless of the generation paradigm—diffusion-based [13], autoregressive [27], or otherwise—there inevitably exists a concluding step that transforms intermediate representations into pixels, serving as the generator’s “final touch.”
We approach this question from the perspective of architecture. To clarify, the generation architecture refers to network components and how they are interconnected. A network component is a functional block composed of multiple neural layers, such as a VAE decoder, DiT transformer, or tokenizer. In contrast, most existing deepfake detection approaches focus on the generation model, defined as a specific set of trained weights (e.g., a pretrained network and its fine-tuned variant are considered different models). Current approaches typically generate training samples from specific generation models and design sophisticated features—such as CLIP embeddings [44], frequency-domain cues [28], multi-scale representations [19], reverse-engineered traces [16], or gradient-based signals [10]—to distinguish synthetic images from real ones. However, as new generators continue to emerge, the generalizability of existing detection approaches becomes a critical concern. When the distribution of testing data shifts significantly, these approaches often require regenerating of synthetic images from new models in order to retrain the detectors. The challenge is further amplified when users fine-tune models on custom data and keep them private.
Figure 1: The proposed taxonomy of generation architectures based on final component.
Since real images are readily available from public datasets for image-related tasks (e.g., ImageNet, MS-COCO), we propose to “contaminate” real images with generation traces using only the generator’s final component, and subsequently train a detector to differentiate them from the original real images. Operating samples through a single component is substantially faster than executing the entire generation pipeline, particularly in diffusion models that require iterative multi-step inference. Moreover, this strategy does not necessitate the generator to be open-sourced; gray-box access to its final component suffices for implementation.
The potential generalizability of the proposed approach is motivated by three key observations: (a) many generators share similar or even identical final components; (b) relying on only a small fraction of the total parameters reduces overfitting to specific generators; (c) the contaminated images preserve identical semantic content to their real counterparts, ensuring that the detector focuses on generation artifacts rather than semantic information specific to the training set.
To comprehensively validate our hypothesis regarding the generalizability of the generator’s final component, we propose a novel taxonomy that categorizes image generators according to their final architectural component and analyze 21 widely used generators under this framework (see Fig.1). Traditionally, image generators are grouped by generation paradigm into three major categories [40]: Generative Adversarial Networks, Diffusion models, and more recently, Autoregressive models. However, many recent generators frequently reuse common components based on their function, such as the VAE decoder, which decodes images from latent space.
From our taxonomy, we identify three representative final components—VAE decoder, super-resolution diffuser, and VQ de-tokenizer—and employ them to imprint traces onto real images as part of our approach. We observe that these components correspond to distinct trace spaces. We conduct a series of experiments to investigate: (a) whether the final component alone is sufficient to detect images generated by its corresponding full architecture; (b) whether a detector trained on one component generalizes to images generated by architectures from other categories; and (c) whether the detector remains effective on fine-tuned variants of the same component. Finally, we select 100 samples from each of the three components using K-medoids clustering, combining the 300 samples to fine-tune a detector built upon a pretrained DINOv3 [32] backbone with an appended fully-connected layer. The resulting detector achieves strong zero-shot performance across multiple challenging benchmarks.
The contributions of this study are as follows:
  • We propose examining the final component of image generation architectures as a source of identifiable traces, enabling the generalizable detection.
  • We introduce a novel taxonomy of text-to-image generators based on their final architectural component, which provides insight into generalization between generators.
  • By imprinting traces onto real images and training a fine-grained, region-based detector on them, our model achieves state-of-the-art zero-shot performance across generator categories and challenging benchmarks.
2 Related Work
2.1 Architecture attribution
For the deepfake detection task that this study focuses on, exploration from the perspective of generative architectures remains limited. In the related task of synthesized image attribution, which aims to identify the generator responsible for a given image, few studies have investigated attribution at the architectural level. Yang et al. [43] claimed to be the first to explore architecture attribution. They assessed GAN architectures and found that architectures leave globally consistent fingerprints, whereas model weights produce regional traces. [41] and [42] reveal that similar diffusion architectures produce similar patterns in frequency-domain features and achieve similar detection accuracies. [1] proposed a Siamese network to verify whether two images originate from the same generation architecture, while [31] applied one-class classification for open-set architecture scenarios.
2.2 Last step inversion
Few studies have examined the final step of the generation process. Laszkiewicz et al.[17] extracted features of generated images by approximating the inputs just before the final layer. The last step they focused on is a single-layer, e.g., activations, fully connected layer, or attention heads. In contrast, our study targets middle grained functional components. [39] proposed a reverse-engineering approach to map the images back to latent features. To effective inverting the decoder, they used the output from the corresponding encoder as the initialization. While related to our sample construction method for the VAE category, their work focused on model attribution and was limited to latent-based generators.
2.3 Zero-shot deepfake detection
In the zero-shot setting, the detection model is trained on samples from certain generators and applied directly to test images without fine-tuning or any prior knowledge of the generating models. Existing zero-shot methods can be broadly grouped into three categories. Signal-level fingerprint construction extracting low-level artifacts as detection cues, such as image gradients (LGrad [10]), pixel neighborhood differences (NPR [18]), or generator noise patterns (DIRE [16]); CLIP-based approaches relying solely on CLIP image features extracted from different transformer layers (UnivFD [46], RINE [23]), or incorporating text guidance using captions or prompts (e.g., “real” vs. “AI-generated”) to steer classification (DeFake [25], C2P-CLIP [44], LASTED [6]); Robust feature design building more resilient features, such as fusing frequency-space features (AIDE [28]) or learning multi-scale features through contrastive learning (CoDE [19]). However, none of these methods explicitly account for the influence of the generator’s architecture or components on the generated images—a gap our approach aims to address.
3 Method
We aim to investigate whether the final component leaves identifiable traces in generated images when all preceding components are omitted, and how well such traces generalize across different generation architectures. The core idea is to use real images to elicit generation signatures and train a binary classifier on them. Let the training set of real images be denoted as $\mathcal{D}=\{\mathbf{x}\}$, where each image $\mathbf{x}\in\mathbb{R}^{h\times w\times 3}$ is labeled as 1 (indicating a real image). Let $\varphi^{*}(\cdot):\mathbb{S}^{d}\rightarrow\mathbb{R}^{h\times w\times 3}$ represent the final component of a well-trained generative model, where $\mathbb{S}^{d}$ is an internal space. To simplify notation, we do not distinguish the model’s parameter set from its corresponding mapping function, e.g., $\varphi^{*}$ denotes both the mapping function implemented by a network module and its learnable parameters. The objective function of the proposed AI-generated image detection is the standard cross-entropy on real images and reconstructed counterparts, as in Eq. 1:
$\mathcal{L}(\theta;\mathcal{E}):=-E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta(\mathbf{x})+log(1-\theta(\varphi^{*}(\mathcal{E}(\mathbf{x})))]$ (1)
where $\mathcal{E}(\cdot):\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{S}^{d}$ is a function that maps RGB image into internal space. In generative models, $\varphi^{*}$ is optimized from the objective of ELBO:
$(\varphi^{*},\mathcal{E}^{*})=$ $\arg\min_{\varphi,\mathcal{E}}\Bigl(-E_{\mathbf{z}\sim q_{\mathcal{E}}(\mathbf{z}|\mathbf{x})}[\log(p_{\varphi}(\mathbf{x}|\mathbf{z}))]+\bigr$.
$\bigl.D_{KL}(q_{\mathcal{E}}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))\Bigr)$ (2)
Given a test image $\mathbf{x}_{t}$, the classifier trained with Eq. 1 outputs $\theta(\mathbf{x}_{t})$, indicating whetherr $\mathbf{x}_{t}$ is real ($1$) or AI-generated (0).
Property 1 If $\mathcal{E}^{*}$ is used as the mapping function $\mathcal{E}$ in Eq.1, then the expected detection error for images generated via $\varphi^{*}$ is not greater than that obtained using any other mapping function $\mathcal{E}$, i.e., $E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta_{\mathcal{E}^{*}}(\mathbf{x})]-E_{\mathbf{z}\sim q_{\mathcal{E}^{*}}(\mathbf{z}|\mathbf{x})}[log(1-\theta_{\mathcal{E}^{*}}(\varphi^{*}(\mathbf{z}))]\leq-E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta_{\mathcal{E}}(\mathbf{x})]-E_{\mathbf{z}\sim q_{\mathcal{E}^{*}}(\mathbf{z}|\mathbf{x})}[log(1-\theta_{\mathcal{E}}(\varphi^{*}(\mathbf{z}))]$, where $\theta_{\mathcal{E}}=\arg\min\mathcal{L}(\cdot;\mathcal{E})$ and $\theta_{\mathcal{E}^{*}}=\arg\min\mathcal{L}(\cdot;\mathcal{E}^{*})$.
This is an idealized property concerning global minimizers, yet it suggests that using the optimal mapping function $\mathcal{E}^{*}$, which was jointly trained with $\varphi^{*}$ in the original generator, results in a lower expected detection error.
[…]
$\hat{\nabla}_{\theta}^{\text{paired}}=-\frac{1}{B}\sum_{\mathbf{x}\in\mathcal{B}_{t}}\Big[\nabla_{\theta}\log\theta(\mathbf{x})+\nabla_{\theta}\log\big(1-\theta(\varphi(\mathcal{E}(\mathbf{x})))\big)\Big]$
$\hat{\nabla}_{\theta}^{\text{indep}}=-\frac{1}{B}\sum_{\mathbf{x}\in\mathcal{B}_{t}}\nabla_{\theta}\log\theta(\mathbf{x})-\frac{1}{B}\sum_{\mathbf{x}^{\prime}\in\mathcal{B}^{\prime}_{t}}\nabla_{\theta}\log\big(1-\theta(\mathbf{x}^{\prime})\big),$
where $\mathcal{B}_{t}\subseteq\mathcal{D}$ and $\mathcal{B}^{\prime}_{t}\subseteq\mathcal{D}^{\prime}$ are mini-batches of size $B$.
A theoretical and empirical discussion is provided in Appendix. This property suggests that independently shuffling real and constructed samples reduces the variance of the stochastic gradient, which in turn leads to more stable and faster convergence in SGD [29, 3]. Therefore, it is advantageous to construct the negative sample set $\mathcal{D}^{\prime}$ offline and sample from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ independently during SGD training.
Figure 2: The proposed framework for detecting AI-generated images
Fig.2 presents an overview of our method. We next detail the two key components of the objective function Eq.1: the fake-sample construction $\varphi\circ\mathcal{E}(\cdot)$ and the detector $\theta$. Before delving into these, we examine the final components employed by popular text-to-image generators and propose a taxonomy to systematically analyze their architectural design choices.
3.1 A Taxonomy of Generation Architectures based on Final Component
The design principles of our taxonomy, grounded in the last component $\varphi(\mathbf{z}):\mathbb{S}^{d}\rightarrow\mathbb{R}^{h\times w\times 3}$, are as follows: (1) Position—Last: The output of $\varphi(\mathbf{z})$ must be an image in RGB space $\mathbb{R}^{h\times w\times 3}$. We define the taxonomy categories based on the input $\mathbf{z}$ and its input space $\mathbb{S}^{d}$. (2) Granularity—Middle-level: Deep network units can be considered at four levels of granularity. The finest level refers to individual layers, such as activation or pooling layers. The second level refers to blocks composed of multiple layers that serve highly detailed functions, such as upsampler in U-Net or transformer heads. The coarsest level is the entire model. “Component” lies in the third level, representing a functional module that operates independently, such as VAE decoders or VQ de-tokenizer. This choice is motivated by two factors: many generators release pre-trained weights for such components separately, making them more accessible without the full pipeline; and fine-grained units, like individual layers or small blocks, leave traces that are too subtle or indistinct for reliable detection. (3) Perspective—Implementation. The taxonomy is based on functional implementation rather than the high-level role of the component. For example, MAR [20] performs autoregressive prediction over masked tokens but ends with a diffusion denoiser, so it is categorized as “denoiser” rather than “de-tokenizer” like other autoregressive generators. DeepFloyd IF [9] performs super-resolution in latent space and ends with a VAE decoder, placing it in the decoder category rather than denoiser like other super-resolution generators.
As illustrated in Fig.1, we define four primary categories based on the input space $\mathbb{S}^{d}$, and further subdivide them according to what $\mathbf{z}$ represents:
  • VAE.decoder. In this category, the final component operates in a continuous latent space, i.e., $\mathbb{S}^{d}:=\mathbb{R}^{d}$, where $d\ll h\times w\times 3$. It decodes a $d$-dimensional feature map or vector into the final image. As many modern generators target high-resolution outputs (e.g., $2048\times 2048$), they often perform generation in a low-dimensional latent space for computational efficiency, relying on encoder–decoder structures for dimensionality reduction and expansion. There are three typical functional contexts in which $\mathbf{z}$ resides: (1) Latent diffusion, where diffusion models operate directly in the latent space, as in the Stable Diffusion family [34] and FLUX [5]. As illustrated in Fig.1(a), in this case, $\mathbf{z}$ represents the denoised latent sample predicted by the diffusion model. (2) Enhancement, where the latent representation from the diffusion model is further refined—e.g., via GAN-based distillation—to produce a more precise $\mathbf{z}$ before decoding, as in HiDream [4]. (3) Latent super-resolution, where a low-resolution image is first generated and then progressively upscaled through latent-space super-resolution stages. As shown in Fig.1(b), the low-resolution image is noised, resized to the target size, and encoded into a latent space for diffusion processing. The resulting upscaled latent, $\mathbf{z}$, is then decoded to produce the final high-resolution image. Examples include CogView 4 [45] and DeepFloyd IF [9], which employ the Stable Diffusion 4× Upscaler [30] as the final super-resolution stage.
    Algorithm 1 Negative sample construction from VAE_1.1
    Input: A real image set $\mathcal{D}=\{\mathbf{x}\}$
    Output: The constructed image set $\{\hat{\mathbf{x}}\}$
    1: vae=AutoencoderKL.from_pretrained(“sd-2-1”, subfolder=“vae”)
    2:for each $\mathbf{x}$ in $\mathcal{D}$ do
    3:  $\mathbf{z}$ = vae.encode($\mathbf{x}$)
    4:  $\hat{\mathbf{x}}$ = vae.decode($\mathbf{z}$)
    5:end for
    6:return $\{\hat{\mathbf{x}}\}$
    Algorithm 2 Negative sample construction from VQ_2.2
    Input: A real image set $\mathcal{D}=\{\mathbf{x}\}$
    Output: The constructed image set $\{\hat{\mathbf{x}}\}$
    1: vq=AutoModelForCausalLM.from_pretrained(“janus”).vq_model
    2:for each $\mathbf{x}$ in $\mathcal{D}$ do
    3:  $\mathbf{z}$ = vq.quantize(vq.encode($\mathbf{x}$))
    4:  $\hat{\mathbf{x}}$ = vq.decode($\mathbf{z}$)
    5:end for
    6:return $\{\hat{\mathbf{x}}\}$
    Algorithm 3 Negative sample construction from Super-resolution_3.3
    Input: A real image set $\mathcal{D}=\{\mathbf{x}\}$
    Output: The constructed image set $\{\hat{\mathbf{x}}\}$
    1: denoiser=PixelFlow(“model.pt”)
    2:for each $\mathbf{x}$ in $\mathcal{D}$ do
    3:  $\mathbf{z}=\hat{\mathbf{x}}_{T}=\alpha\mathbf{x}+\beta\epsilon$, $\epsilon\sim N(\mathbf{0},\mathbf{I})$
    4:  for t=T1t=T-1 to 0 do
    5:   $\tilde{\epsilon}$=denoiser($\hat{\mathbf{x}}_{t+1}$)
    6:   $\hat{\mathbf{x}}_{t}=$ step($\hat{\mathbf{x}}_{t+1}$, $\tilde{\epsilon}$))
    7:  end for
    8:end for
    9:return $\{\hat{\mathbf{x}}_{0}\}$
  • VQ.de-tokenizer. The final component operates in discrete token space, i.e., $\mathbb{S}^{d}:=\mathbb{Q}^{p}$, where $\mathbb{Q}$ is a learned codebook and $p$ denotes the number of spatial tokens (i.e., divided image patches) in the quantized representation. There are three common contexts for $\mathbf{z}$: (1) Next-token auto-regression, where tokens are predicted sequentially and the VQ decoder transforms them back into image pixels, as in Emu 3 [38], JanusPro [8], and LlamaGen [35]. (2) Next-scale autoregression, where the generator starts from coarse-scale tokens (e.g., a $1\times 1$ token map) and progressively predicts higher-resolution token maps conditioned on all previous ones, as in VAR [36] and Infinity [12]. (3) Latent diffusion hybrid, which combines global latent diffusion for coarse image structure with autoregressive token prediction for sequentially refining discrete latent components, e.g., Kandinsky [2].
  • Diffusion.denoiser. In this category, the final step is denoising, applied over three typical input spaces. (1) Continuous token: $\mathbb{S}^{d}:=\mathbb{R}^{p}$. The image is divided into $p$ patches, with tokens represented as continuous embeddings rather than discrete codes. MAR [20] exemplifies this approach, as illustrated in Fig.1(d). (2) Low-resolution image: $\mathbb{S}^{d}:=\mathbb{R}^{h_{l}\times w_{l}\times 3}$, where $h_{l}<h,w_{l}<w$. In this setting, the generator builds images progressively from low to high resolution through multiple upscaling stages, e.g., $64\times 64\rightarrow 256\times 256\rightarrow 1024\times 1024$. The final step adds noise to the previous resolution’s image, resizes it to the target size, and applies diffusion denoising. This process is similar to VAE.decoder–latent super-resolution but operates directly in pixel space (see Fig.1(e)). Two training strategies are common: cascade super-resolution, where each adjacent low-to-high resolution stage is trained separately (e.g., DALL·E 3 [26] and GLIDE [24]), and end-to-end super-resolution, where all stages are trained jointly (e.g., PixelFlow [7]). (3) Same-size pixels: $\mathbb{S}^{d}:=\mathbb{R}^{h\times w\times 3}$. The input $\mathbf{z}$ has the same spatial resolution as the target image. It can either represent a low-quality image to be refined by the diffusion denoiser, as in ReflectionFlow [47], or an initial Gaussian noise sample used in conventional diffusion models such as DDPM [13] and DDIM [33] (Fig.1(f)).
  • One-stop.generator. In this setting, $\mathbb{S}^{d}:=\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. The generator directly maps a sampled noise vector to the final image, as shown in Fig.1(g). This category includes GAN-based approaches such as GigaGAN [14], which produce high-resolution images through large-scale adversarial training, and pixel-wise auto-regressive models like JetFormer [37], which generate images sequentially in an end-to-end fashion.
The categorization of these representative generators in our proposed taxonomy is based on their official source code or released documentation, with detailed references provided in the Appendix.
3.2 Fake sample construction — $\varphi\circ$ $\mathcal{E}$
3.2.1 Construction from the three representative components
To investigate the generalization of final component $\varphi$ in generated trace detection (Eq. 1), we identify three representative final components from open-source generators in our taxonomy as shown in Fig.1: Stable Diffusion 2.1 (VAE.decoder-1.1:Latent diffusion), JanusPro (VQ.de-tokenizer-2.2:Next-token auto-regression), and PixelFlow (Diffusion.denoiser-3.3:End-to-end super-resolution). For brevity, we refer to these specific components as VAE_1.1, VQ_2.2, and Super-resolution_3.3 respectively in the following sections. As discussed in Property 1 (Sec. 3), the encoder $\mathcal{E}^{*}$, which is pretrained jointly with $\varphi^{*}$, is used to approximate the latent input $\mathbf{z}$.
Algorithm 1 presents the pseudocode for constructing $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ using the VAE_1.1. The implementation requires only a few lines of code and is significantly more efficient than executing the full generative pipeline. For VQ_2.2, the de-tokenizer acts as $\varphi^{*}$ and the image tokenizer as $\mathcal{E}^{*}$ to approximate image tokens $\mathbf{z}$, as outlined in Algorithm 2. In Super-resolution_3.3, the input $\mathbf{z}$ is a noisy version of the real image, $\alpha\mathbf{x}+\beta\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,I)$. The noise level can be adjusted by selecting the starting timestep $t$ for denoising (Algorithm 3).
Figure 3: Performance comparison of our model trained on the final component of SD2
Figure 4: Performance comparison of our model trained on the final component of HiDream
Figure 5: Feature visualization of the three final components.
3.2.2 Sparse Sampling from Combined Constructions
The three final components differ substantially in both functionality and implementation. Consequently, the constructed samples $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ from each reflect distinct generative traces. We further investigate whether combining a small number of representative fake samples from each component can better capture the overall trace space and improve generalization. Algorithm 4 outlines the procedure, applying K-Medoids clustering selects $k$ representative samples from the constructed outputs of each component, as described in Section 3.2.1. The selection is performed in a reduced-dimensional feature space. Specifically, the generated samples are first encoded using a pretrained DINOv3 backbone, from which 2,048-dimensional feature vectors are extracted from the final layer. K-Medoids selects $k$ samples such that most other samples are close to one of them in Euclidean distance. The selected subsets from all three components are combined to form the final training set of $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ for detector training.
3.3 Detection model — $\theta$
Algorithm 4 Negative sample selection
Input: Constructed image sets $\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\mathcal{D}_{3}=\{\hat{\mathbf{x}}\}_{VAE}\cup$ $\{\hat{\mathbf{x}}\}_{VQ}\cup\{\hat{\mathbf{x}}\}_{sup}$, the number of samples to be selected $k$
Output: The sparse constructed image set $\{\hat{\mathbf{x}}\}$
1: dinov3 = torch.hub.load(“dinov3_vith16plus”)
2:for $i=1$ to 33 do
3:  features.append(dinov3($\hat{\mathbf{x}}$)) for all $\hat{\mathbf{x}}\in\mathcal{D}_{i}$
4:  dist = sklearn.metrics.euclidean_distances(features)
5:  $\mathcal{M}_{i}$= sklearn.kmedoids.fasterpam(dist, $k$) //a set of $k$ samples of $\hat{\mathbf{x}}$
6:end for
7:return $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$
Based on Property 2 discussed in Section 3, instead of loading $\mathbf{x}$ and computing $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ at each training step, we directly input the real image set (positive samples) $\{\mathbf{x}\}$ and the constructed image set $\{\hat{\mathbf{x}}\}$ generated by Algorithms 14, with standard shuffling applied to both sets during SGD training. Since the detector $\theta$ aims to distinguish subtle differences between real and reconstructed images, and object detection tasks inherently emphasize multi-scale spatial features, we adopt DINOv3—a foundation model pretrained for object detection—as the backbone to better capture fine-grained artifacts. A single fully connected layer is appended to the backbone, and all layers are fine-tuned during training. As our focus is on analyzing the effect of the final generative component, we employ a pretrained visual backbone with minimal modifications rather than designing a complex detector.
4 Experiments
4.1 Model implementation & baselines
The real image set used for constructing the training dataset is MS-COCO 2014 [21], which contains 118,000 images depicting diverse indoor and outdoor scenes with resolutions ranging from 384 to 880 pixels in width (height). The constructed training and corresponding test sets are described separately for each of the following experimental questions, with all pretrained models referenced in the Appendix.
For our AI-generated image detector, we use the pretrained backbone of DINOv3111https://huggingface.co/facebook/dinov3-vith16plus-pretrain-lvd1689m with an additional fully connected layer for binary classification. All layers are fine-tuned using standard cross-entropy loss with a learning rate of 5e-7. This implementation serves as the detector for all subsequent analyses.
We compare our method against eight baseline approaches: AIDE [28] (ICLR 2025), C2P-CLIP [44] (AAAI 2025), CoDE [19] (ECCV 2024), RINE [23] (ECCV 2024), FatFormer [22] (CVPR 2024), NPR [18] (CVPR 2024), LGrad [10] (CVPR 2023), and DIRE [16] (ICCV 2023). As none of these methods report results on our testing sets, we re-evaluate all of them using their publicly released pretrained models and default hyperparameters. The source code links, parameters, and models used are detailed in the Appendix. Performance is assessed using two standard metrics: classification accuracy and average precision (AP).
4.2 Results & discussions
Table 2.Detection results across different category of generators.
vae.decoder based diffusion.denoiser.super-resolution based vq.decoder based
SD1.3 SD1.4 SD XL Flux SD 3.5 DALL·E 2 DALL·E 3 Glide Emu3 LlamaGen
Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP
DIRE (ICCV2023) 48.75 0.43 48.65 0.43 50.00 0.59 61.66 0.73 55.65 0.59 56.45 0.67 50.35 0.54 49.50 0.49 53.65 0.57 54.05 0.59
LGrad (CVPR2023) 54.45 0.46 55.85 0.48 52.60 0.48 81.91 0.92 77.29 0.87 55.65 0.56 47.65 0.40 47.45 0.52 54.80 0.48 60.70 0.63
NPR (CVPR2024) 50.70 0.51 50.60 0.52 52.45 0.56 90.00 0.92 88.04 0.90 49.25 0.50 50.35 0.51 48.80 0.51 51.70 0.53 52.25 0.55
FatFormer (CVPR2024) 52.97 0.54 53.72 0.54 66.67 0.75 55.32 0.90 75.86 0.95 54.12 0.56 35.88 0.34 67.77 0.76 63.72 0.66 77.21 0.87
RINE (ECCV2024) 87.60 0.96 87.15 0.96 74.60 0.86 51.70 0.87 96.49 0.99 85.90 0.94 42.35 0.31 63.95 0.72 62.10 0.70 90.65 0.98
CoDE (ECCV2024) 97.75 0.97 97.30 0.97 75.20 0.75 71.09 0.72 76.35 0.75 54.60 0.54 73.15 0.73 80.75 0.80 77.45 0.77 90.50 0.90
C2P-CLIP (AAAI2025) 75.15 0.93 76.70 0.93 77.70 0.95 53.96 0.97 80.88 0.99 55.60 0.77 63.20 0.90 86.65 0.97 68.70 0.91 83.80 0.98
AIDE (ICLR2025) 57.60 0.60 56.70 0.60 62.70 0.64 98.50 0.99 96.45 0.99 59.75 0.43 14.15 0.31 59.75 0.77 62.50 0.57 63.25 0.85
Ours_vae_1.1 99.70 1.0 99.70 1.0 99.70 1.0 99.17 .9998 99.96 1.0 99.60 1.0 99.70 .9999 98.85 .9992 99.70 .9999 99.70 .9999
Ours_vq_2.2 99.85 1.0 99.85 1.0 99.85 1.0 99.19 .9997 99.82 .9999 99.80 1.0 99.85 .9999 98.75 .9984 99.85 .9999 99.85 1.0
Ours_super-res_3.3 99.90 1.0 99.80 1.0 99.90 1.0 99.89 1.0 99.86 1.0 99.35 .9991 99.85 1.0 98.60 .9985 99.90 1.0 99.90 1.0
Ours_sparse 99.15 .9990 99.15 .9990 99.15 .9988 98.95 .9828 98.98 .9958 99.80 .9987 99.15 .9991 98.05 .9947 99.15 .9990 99.15 .9982
In experiments, we aim to address the following questions:
Q1: Are traces from the final component alone sufficient to detect images generated by its full pipeline?
To verify this, we train two detectors on samples from the VAE decoders of SD2 and HiDream-I1 and evaluate them on images produced by their respective full generation pipelines.
Dataset: (1) SD2. The training samples comprise real MS-COCO images and corresponding fake images constructed via Algorithm 1 using Stable Diffusion 2.1. The test set includes 1,000 real images and 1,000 SD2-generated images from Synthbuster [^2] https://zenodo.org/records/10066460. (2) HiDream-I1. The fake training samples are constructed in same procedure but using the HiDream-I1 model, and the test set is generated from the same 1,000 prompts as Synthbuster.
Fig.s 5 and 5 show that both detectors perform strongly. For SD2, baselines vary widely. CoDE and C2P-CLIP achieve relatively high results because their training sets include SD2-generated images, making this an in-distribution evaluation. In contrast, HiDream-I1 evaluation is fully zero-shot. For the AP metric, our detector achieves nearly perfect scores, with logits for real samples approaching 1 and those for fake samples approaching 0. These results demonstrate that training with the generator’s final component generalizes effectively to the generator’s full outputs.
Q2: Are traces from the final component sufficient to detect images from full pipelines across categories?
Table 3: Detection results on unknown generators.
Unknown Mixture Wild
Firefly Midjourney FakeBench Reddit Facebook Twitter
Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP
DIRE 50.8 0.55 53.9 0.61 44.38 0.42 57.19 0.62 56.56 0.66 47.74 0.79
LGrad 47.25 0.42 57.7 0.53 44.01 0.69 57.80 0.65 66.88 0.77 40.86 0.61
NPR 47.30 0.47 52.35 0.55 52.37 0.58 63.80 0.67 74.06 0.84 48.46 0.68
FatFormer 60.12 0.62 44.43 0.42 71.2 0.81 69.47 0.81 54.83 0.75 42.79 0.77
RINE 92.10 0.99 55.45 0.59 74.77 0.97 71.20 0.81 51.56 0.55 55.93 0.62
CoDE 60.65 0.60 76.90 0.77 74.75 0.74 65.8 0.61 72.19 0.67 68.41 0.78
C2P-CLIP 59.45 0.82 52.85 0.69 74.77 0.91 68.40 0.79 54.37 0.71 47.27 0.75
AIDE 14.40 0.32 53.45 0.40 74.78 0.87 70.87 0.83 50.00 0.56 52.78 0.59
Ours_vae_1.1 99.70 1.0 99.70 .9999 98.83 .9997 96.87 .9968 97.81 .9899 98.34 .9993
Ours_vq_2.2 99.70 .9999 99.45 .9996 98.10 .9982 95.40 .9980 95.31 .9928 98.34 .9986
Ours_sup-res 98.50 .9996 99.70 1.0 94.38 .9978 98.67 .9964 97.19 .9946 97.39 .9998
Ours_sparse 98.00 .9973 99.10 .9990 97.90 .9991 97.60 .9951 98.44 .9944 99.31 1.0
Table 4: Results on finetuned generators.
Satellite image generation
amuse- park car dealership electric station stadium
Acc AP Acc AP Acc AP Acc AP
50.49 0.86 50.19 0.84 50.40 0.84 50.58 0.94
48.85 0.36 50.00 0.42 50.15 0.47 50.02 0.43
46.89 0.43 65.52 0.84 65.65 0.93 56.16 0.85
51.68 0.70 50.88 0.83 52.81 0.79 51.66 0.84
69.07 0.83 51.97 0.50 64.19 0.74 68.16 0.78
86.66 0.82 88.40 0.84 84.50 0.80 78.37 0.73
58.00 0.76 76.60 0.96 71.71 0.88 58.74 0.94
52.90 0.58 61.71 0.85 61.11 0.76 52.32 0.72
99.33 1.0 96.43 .9996 96.08 .9987 96.50 .9997
96.77 1.0 87.52 .9998 90.99 .9989 98.89 .9999
99.67 .9997 97.94 .9962 96.50 .9917 97.34 .9943
99.61 .9999 98.91 .9994 98.55 .9982 99.59 .9998
We train detectors on the three representative final components (Section 3.2.1) and an additional detector on the sparse image set (Section 3.2.2), then evaluate them on images from different categories in our taxonomy.
Dataset: Training sets. For the VAE_1.1, VQ_2.2, and Super-Resolution_3.3 components, real images are from MS-COCO, and fake samples are constructed following Algorithms 13. These three types of final components exhibit distinct generation traces, visualized in Fig. 5. Features are extracted from the third layer of the pretrained DINOv3 backbone and project into three dimensions. The resulting distributions appear mixed yet remain visually separable. The sparse training set is selected using Algorithm 4, with $k=100$. As shown in Fig. 6, the selected data points (highlighted in red) span the entire space. Testing set. The testing set consists of image subsets from the Synthbuster benchmark, generated by SD1.3, SD1.4, SD XL, DALL·E 2, DALL·E 3, and Glide. We categorize these subsets according to our taxonomy, as summarized in Table 6. For the VQ decoder category, we generate fake images from two autoregressive generators, Emu3 and LlamGen, using the 1,000 Synthbuster prompts. For emerging generators Flux-dev and SD3.5, we produce 10,000 fake images each. As paired real images are unavailable, we use 9,600 real images from the PASCAL VOC training split [11] for evaluation.
Figure 6: Sparse sample selection.
Table 6 reports detection results across generator categories. Each row in the last group corresponds to a detector trained on one of the four training sets—VAE_1.1, VQ_2.2, Super-resolution_3.3, or the sparse set. Our approach performs strongly within the same category and generalizes well across categories. For example, on DALL·E 2—where most baselines are near random (e.g., NPR: 49.20%, LGrad: 55.65%, FatFormer: 54.12%)—only CLIP-based models like RINE (85.90%) and AIDE (59.75%) show moderate gains, benefiting from DALL·E 2’s CLIP-based design. Despite not relying on CLIP features, our method achieves substantially higher accuracy, highlighting robust generalization to unseen generator architectures.
Notably, the detector trained on the sparse set, using only 100 fake samples from each of the three components (300 fake and 300 real images in total), achieves performance comparable to our three models trained on individual components with the full MS-COCO dataset. This small training set provides further evidence of the generalization capability of our approach and offers a practical advantage, as generating large-scale training images is costly.
Q3: Are traces from the final component sufficient to detect images from wild or unknown generators?
The detectors trained for Q2 are directly evaluated on testing sets for this question.
Dataset: Testing set. (1) Unknown generators — Firefly and Midjourney test sets from Synthbuster. Both are commercial generators lacking open-source code or public documentation of their architectures or components. (2) Mixed generators—FakeBench  [^3] https://github.com/Yixuan423/FakeBench benchmark contains 3,000 real images and 3,000 fake images generated by multiple models. (3) In-the-wild generation—WildRF  [^4] https://vision.huji.ac.il/ladeda/ comprises three subsets: Reddit, Facebook, and Twitter, each containing real and synthetic images sourced from social media platforms with unknown origins.
Table 4 reports the results. Baseline methods exhibit highly inconsistent results across benchmarks. For instance, RINE performs well on Firefly (92.10% accuracy, 0.99 AP) but drops sharply on Midjourney (55.45%) and Facebook (51.56%), while AIDE misclassifies most samples—achieving only 14.40% accuracy. This instability highlights the challenge of real-world generalization, particularly in true zero-shot scenarios where no prior information about the generators is available. In contrast, our detector consistently surpasses all baselines and maintains robust, stable performance across all subsets, demonstrating that leveraging traces from the final component yields strong generalization to diverse architectures and challenging real-world conditions.
Q4: Does detection performance on a full generator remain valid after fine-tuning the generator?
The detectors trained for Q2 are directly evaluated on testing sets for this question.
Dataset: Testing set — SatelliteDiffusion[15] dataset is generated using a domain-specific model fine-tuned from SD2, conditioned on satellite metadata. Real images are sourced from high-resolution satellite imagery, while synthetic counterparts are produced using textual descriptions of the real samples. Four subsets: amusement park, car dealership, electric substation, and stadium are tested, each including around 2,000 real and 2,000 generated images.
Table 4 presents the comparison. Our models attain results closely matching those reported for the original SD2 evaluations (Fig. 5, Table 6), demonstrating that detection remains effective for generators fine-tuned on domain-specific images and conditions.
Figure 7: Performances of our detectors with different backbones
Q5: How does the classification backbone affect detection performance? To evaluate the impact of the backbone on our detector, we train a model using a pretrained DINOv2 backbone on the VAE training set used in the previous experiments. We compare the performance of DINOv2 and DINOv3 on the testing sets from Q3 and Q4 (Fig. 7). The results show that the newer backbone provides a significant improvement in accuracy. For AP, since the margin between the logits of positive and negative images is sufficiently large with DINOv2, using the newer backbone has little effect.
5 Conclusion
This study demonstrates that leveraging the final components of modern image generators offers a powerful means to enhance deepfake detection generalization. By contaminating real images through these components and training on minimal representative samples, the proposed detector achieves superior performance across diverse unseen generators, highlighting the value of insights for robust detection.
Supplementary Material
6 Discussion on Property 2
For convenience, the paired gradient $\hat{\nabla}_{\theta}^{\text{paired}}$ can be rewritten as:
$\hat{\nabla}_{\theta}^{\text{paired}}=-\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\Big[\nabla_{\theta}\log\theta(\mathbf{x}_{i})+\nabla_{\theta}\log(1-\theta(\varphi(\mathcal{E}(\mathbf{x}_{i}))))\Big].$
Define
$g_{i}=\nabla_{\theta}\log\theta(\mathbf{x}_{i}),\qquad g_{i}^{\prime}=\nabla_{\theta}\log(1-\theta(\varphi(\mathcal{E}(\mathbf{x}_{i})))),$
and let
$G=\{g_{i}\}_{i=1}^{|\mathcal{B}_{t}|},\qquad G^{\prime}=\{g_{i}^{\prime}\}_{i=1}^{|\mathcal{B}_{t}|}.$
Under this notation, $\hat{\nabla}_{\theta}^{\text{paired}}$ becomes
$\hat{\nabla}_{\theta}^{\text{paired}}=-\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\big[g_{i}+g_{i}^{\prime}\big].$
Since the elements of $\mathcal{B}_{t}$ are sampled independently, each $\mathbf{x}_{i}$ is independent of all other samples in $\mathcal{B}_{t}$. Consequently, each gradient $g_{i}$ is independent of all other elements in $G$ except itself, and the same holds for $g_{i}^{\prime}$ within $G^{\prime}$. Furthermore, $g_{i}$ is independent of all elements of $G^{\prime}$ except for its paired counterpart $g_{i}^{\prime}$, and vice versa.
Using the fact that independent random variables have zero covariance, the variance of $\hat{\nabla}_{\theta}^{\text{paired}}$ is:
$\text{Var}(\hat{\nabla}_{\theta}^{\text{paired}})$ $=\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\Big[\text{Var}(g_{i})+\text{Var}(g_{i}^{\prime})+2\text{Cov}(g_{i},g_{i}^{\prime})\Big]$
$=\frac{1}{|\mathcal{B}_{t}|}\Big[\text{Var}(G)+\text{Var}(G^{\prime})+\text{Cov}(G,G^{\prime})\Big],$
where
$\text{Var}(G)=\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Var}(g_{i}),\qquad\text{Var}(G^{\prime})=\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Var}(g_{i}^{\prime}),$
$\text{Cov}(G,G^{\prime})=2\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Cov}(g_{i},g_{i}^{\prime}).$
In contrast, the variance under independent sampling is
$\text{Var}(\hat{\nabla}_{\theta}^{\text{indep}})=\frac{1}{|\mathcal{B}_{t}|}\text{Var}(G)+\frac{1}{|\mathcal{B}_{t}|}\text{Var}(G^{\prime}).$
The only difference between the two variances is the covariance term $\text{Cov}(G,G^{\prime})$, which aggregates the individual covariances between $g_{i}$ and $g_{i}^{\prime}$. Our key observation is that $\mathbf{x}_{i}$ and its transformed version $\mathbf{x}_{i}^{\prime}=\varphi(\mathcal{E}(\mathbf{x}_{i}))$ are highly similar, and thus their gradients $g_{i}$ and $g_{i}^{\prime}$ tend to be positively correlated. Therefore, $\text{Cov}(g_{i},g_{i}^{\prime})>0$, implying $\text{Cov}(G,G^{\prime})>0$. As a result,
$\text{Var}(\hat{\nabla}_{\theta}^{\text{paired}})>\text{Var}(\hat{\nabla}_{\theta}^{\text{indep}}).$
This confirms our claim that the gradient estimator achieves lower variance when mini-batches are sampled independently from $\mathcal{D}$ and $\mathcal{D}^{\prime}$.
We empirically compare pairwise training with independent training. Using samples generated by VAE_1.1, we train detectors with identical model architectures, hyperparameters, and training data. The only difference lies in how mini-batches are constructed: pairwise training loads matched pairs ($\mathbf{x}$, $\mathbf{x}^{\prime}$) into the same mini-batch, whereas independent training samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$ independently and loads them separately, while ensuring that each mini-batch contains the same number of positive samples $\mathbf{x}$ and negative samples $\mathbf{x}^{\prime}$. Table A reports the results. We copy the baseline numbers and Ours_VAE_1.1 results from Table 2 of the main paper, along with the additional row for pairwise training. The average accuracy of pairwise training is 97.22%, compared to 99.22% for independent training—a 2% gap. This empirical observation suggests that pairwise training does not provide an advantage over independent training, consistent with our theoretical analysis above.
Table A. Comparison between pairwise training and independent training.
vae.decoder based diffusion.denoiser.super-resolution based vq.decoder based
SD1.3 SD1.4 SD XL Flux SD 3.5 DALL·E 2 DALL·E 3 Glide Emu3 LlamaGen
Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP
DIRE (ICCV2023) 48.75 0.43 48.65 0.43 50.00 0.59 61.66 0.73 55.65 0.59 56.45 0.67 50.35 0.54 49.50 0.49 53.65 0.57 54.05 0.59
LGrad (CVPR2023) 54.45 0.46 55.85 0.48 52.60 0.48 81.91 0.92 77.29 0.87 55.65 0.56 47.65 0.40 47.45 0.52 54.80 0.48 60.70 0.63
NPR (CVPR2024) 50.70 0.51 50.60 0.52 52.45 0.56 90.00 0.92 88.04 0.90 49.25 0.50 50.35 0.51 48.80 0.51 51.70 0.53 52.25 0.55
FatFormer (CVPR2024) 52.97 0.54 53.72 0.54 66.67 0.75 55.32 0.90 75.86 0.95 54.12 0.56 35.88 0.34 67.77 0.76 63.72 0.66 77.21 0.87
RINE (ECCV2024) 87.60 0.96 87.15 0.96 74.60 0.86 51.70 0.87 96.49 0.99 85.90 0.94 42.35 0.31 63.95 0.72 62.10 0.70 90.65 0.98
CoDE (ECCV2024) 97.75 0.97 97.30 0.97 75.20 0.75 71.09 0.72 76.35 0.75 54.60 0.54 73.15 0.73 80.75 0.80 77.45 0.77 90.50 0.90
C2P-CLIP (AAAI2025) 75.15 0.93 76.70 0.93 77.70 0.95 53.96 0.97 80.88 0.99 55.60 0.77 63.20 0.90 86.65 0.97 68.70 0.91 83.80 0.98
AIDE (ICLR2025) 57.60 0.60 56.70 0.60 62.70 0.64 98.50 0.99 96.45 0.99 59.75 0.43 14.15 0.31 59.75 0.77 62.50 0.57 63.25 0.85
Ours_vae_1.1 99.70 1.0 99.70 1.0 99.70 1.0 99.17 .9998 99.96 1.0 99.60 1.0 99.70 .9999 98.85 .9992 99.70 .9999 99.70 .9999
Ours_vae_1.1_pairwise 98.20 .9999 98.20 .9999 98.20 .9992 98.88 .9542 99.07 .9775 98.10 .9998 98.20 .9970 98.05 .9991 98.20 .9998 98.20 .9998
7 Evidence sources for categorization
Table B summarizes the sources used to determine the final-component category of each generator. For all open-source models, we list the exact commit version, together with the file paths and line numbers that invoke the final component. DALL·E 2 provides only documentation rather than code, and DALL·E 3 has neither detailed documentation nor public source code. Among the listed generators, Flux-1-dev has two widely used official implementations: the Diffusers version operates in latent space and uses a VAE decoder as its final component, while the Black Forest Labs implementation performs generation directly in pixel space. A similar situation appears in DeepFloyd IF: its Stage III super-resolution step has two variants—one based on the Stable Diffusion ×4 latent upsampler, and another implemented in pixel space. Moreover, although MAR is grouped under the “denoiser” category, its final step is a decoder network structurally similar to a U-Net denoiser but without any noise addition or removal.
Table B. Sources for generator categorization.
Name Category Source Evidence Stable Diffusion 3.5 vae.decoder https://github.com/Stability-AI/sd3.5 Commit 106db06
sd3_infer.py : Line 495
image = self.vae_decode(sampled_latent)
Flux 1-dev vae.decoder https://github.com/huggingface/diffusers/tree/main/src/diffusers/pipelines/flux Commit 5e181ed
pipeline_flux.py : Line 1006
image = self.vae.decode(latents, return_dict=False)[0]
diffusion. denoiser https://github.com/black-forest-labs/flux Commit 57ce405
src/flux/sampling.py : Line 351
img = img + (t_prev - t_curr) * pred
HiDream-I1 vae.decoder https://github.com/HiDream-ai/HiDream-I1 Commit 3519729
hi_diffusers/pipelines/hidream_image/ pipeline_hidream_image.py : Line 724
image = self.vae.decode(latents, return_dict=False)[0]
Stable diffusion xl vae.decoder https://github.com/huggingface/diffusers/tree/main/src/diffusers/pipelines/stable_diffusion_xl& Commit 0d1c5b0
pipeline_stable _diffusion_xl.py : Line 1292
image = self.vae.decode(latents, return_dict=False)[0]
CogView4 vae.decoder https://github.com/huggingface/diffusers/blob/main/src/diffusers/pipelines/cogview4 Commit 0d1c5b0
pipeline_cogview4.py : Line 673
image = self.vae.decode(latents, return_dict=False, generator=generator)[0]
DeepFloyd IF vae.decoder https://github.com/deep-floyd/IF/ Commit af64403
deepfloyd_if/modules/stage_III_sd_x4.py : Line 80
available_models = [’stable-diffusion-x4-upscaler’] ...
images = self.model(**metadata).images
diffusers/src/diffusers/pipelines/stable_diffusion Commit a4df8db
pipeline_stable _diffusion_upscale.py : Line 798
image = self.vae.decode(latents / self.vae.config.scaling_factor, return_dict=False)[0]
diffusion. denoiser https://github.com/deep-floyd/IF/ Commit af64403
deepfloyd_if/modules/stage_III.py:Line 26
return super().embeddings_to_image(...)
Commit af64403
deepfloyd_if/modules/base.py : Line 194
sample = diffusion.p_sample_loop(...)
Kandinsky 3.0 vq.de-tokenizer https://github.com/ai-forever/Kandinsky-3 Commit 10db67a
kandinsky3/movq.py: Line 420,421
decoder_input = self.post_quant_conv(quant) decoded = self.decoder(decoder_input, quant)
Emu3 vq.de-tokenizer https://github.com/baaivision/Emu3 Commit 1a43ee6
emu3/mllm/processing_emu3.py : Line 198, 199
doc=self.tokenizer.decode(*args,**kwargs) return self.multimodal_decode(doc)
JanusPro vq.de-tokenizer https://github.com/deepseek-ai/Janus Commit 659982f
janus/models/ vq_model.py: Line 506,507
quant_b = self.quantize.get_codebook_entry() dec = self.decode(quant_b)
LlamaGen vq.de-tokenizer https://github.com/FoundationVision/LlamaGen Commit 5a50452
autoregressive/sample/ sample_t2i.py : Line 121
samples = vq_model.decode_code(index_sample, qzshape)
Infinity vq.de-tokenizer https://github.com/FoundationVision/Infinity Commit 3ab8e0c
infinity/models/infinity.py : Line 636
img = vae.decode(summed_codes.squeeze(-3))
VAR vq.de-tokenizer https://github.com/FoundationVision/VAR Commit a5cf0a1
models/var.py : Line 190
return self.vae_proxy[0].fhat_to_img(f_hat) // self.vae_proxy: Tuple[VQVAE]
MAR decoder https://github.com/LTH14/mar Commit fe1de72
models/mar.py : Line: 225
x = block(x)
DALL·E 2 diffusion. denoiser Document: Hierarchical Text-ConditionalImage Generation with CLIP Latents Page 4 “we train two diffusion upsampler models: one to upsample images from 64×64 to 256×256 resolution, and another to further upsample those to 1024×1024 resolution” GLIDE diffusion. denoiser https://github.com/openai/glide-text2im Commit 1f791b8
glide_text2im/text2im_model.py: Line 167
return super().forward(x, timesteps, **kwargs)
PixelFlow diffusion. denoiser https://github.com/ShoufaChen/PixelFlow Commit eeabc08
pixelflow/pipeline_pixelflow.py : Line 253
latents = self.scheduler.step( model_output=noise_pred,sample=latents)
Reflection- Flow diffusion. denoiser https://github.com/Diffusion-CoT/ReflectionFlow train_flux /sample.py : inference-time refine by black-forest-labs/FLUX.1-dev Flux-1-dev (pixel-space implmentation; see note above) DDPM diffusion. denoiser https://github.com/hojonathanho/diffusion Commit c461299
diffusion_tf/diffusion_utils_2.py : Line 211
self.p_sample(..., x=img_, ...)
DDIM diffusion. denoiser https://github.com/huggingface/diffusers/tree/main/src/diffusers/pipelines/ddim Commit a4df8db
pipeline_ddim.py : Line 152
image = self.scheduler.step(...)
GigaGAN one-stop. generator https://github.com/lucidrains/gigagan-pytorch Commit 0806433
gigagan_pytorch/gigagan_pytorch.py: Line 2167, 2169
model = self.G_ema if self.has_ema_generator else self.G
return model(*args, **kwargs)
JetFormer one-stop. generator https://github.com/google-research/big_vision Commit 9c006bf
big_vision/models/proj/jetformer/ jetformer.py: Line 481
prelogits, decoder_out = self.decoder(x, ...)
8 Baseline models
As none of the baseline methods report results on our testing sets, we re-evaluate all of them using their publicly released pretrained models and default hyperparameters. All baselines compute Accuracy and AP using the same functions from sklearn.metrics: accuracy_score and average_precision_score. The only exception is CoDE, which does not include AP computation in its evaluation code; for this case, we add the same AP calculation used by the other baselines. Below, we provide the links to the testing code and pretrained weights. For each baseline, we use the default hyperparameters specified in the authors’ example scripts within their repositories.
9 Datasets construction
Based on Property 2, we generate the constructed image sets $\mathbf{x}^{\prime}$} offline and save them before training begins. For each of the three representative final components, the corresponding constructed training set is produced using its well-trained, optimized encoder, following the guidance of Property 1. Below, we list the real images and the pretrained $\mathcal{E}^{*}$ and $\varphi^{*}$ used in our experiments:
  • VAE_1.1 training set: The real images are from the full MS-COCO 2014 dataset (118k images). The pretrained VAE used to generate negative samples is the VAE from https://huggingface.co/stabilityai/stable-diffusion-2-1.
  • HiDream training set in Experiment Q1: The pretrained VAE weights are from HiDream-I1-Full https://huggingface.co/HiDream-ai/HiDream-I1-Full.
  • VQ_2.2 training set: The pretrained model used for converting real images is JanusPro-7B https://huggingface.co/deepseek-ai/Janus-Pro-7B. Due to its fixed patch size, Janus only supports a resolution of 378×378. Therefore, we filter out images smaller than this size and center-crop the remaining MS-COCO images to 378×378.
  • Super-resolution_3.3 training set: We use PixelFlow https://huggingface.co/ShoufaChen/PixelFlow-Text2Image/tree/main/model.pt, whose last stage is pretrained to upscale images from 128×128 to 512×512. For real images in MS-COCO, we crop them to 512×512 whenever the resolution allows. We do not downsample and re-upsample the real images; instead, we directly use real images that match the output resolution.
  • Sparse selection set: We run Algorithm 4 on the constructed sample sets from VAE_1.1, VQ_2.2, and Super-resolution_3.3. Once the constructed (negative) samples are selected, their corresponding real images form the positive training set.
For the testing set, we use the full sets from the benchmarks Synthbuster, FakeBench, and WildRF. In Synthbuster, there are 1,000 high-resolution real images from RAISE-1K, along with 1,000 fake images generated by each of nine models: DALL·E2, DALL·E3, Adobe Firefly, MidJourney v5, SD1.3, SD1.4, SD2, SDXL, and Glide. The corresponding columns in Tables 2 and 3 correspond to these subsets. To include more recent generators, we additionally construct the following test sets used in the experiments reported in the main paper:
  • HiDream (Q1), Emu3, and LlamaGen (Q2): Real images are the same as in Synthbuster. Fake images are generated from the same 1,000 prompts used for the original nine generators (listed in prompts.csv in https://zenodo.org/records/10066460/files/synthbuster.zip). Generation parameters and image resolutions follow the defaults in each repository, as listed in Table B.
  • SD3.5 and Flux-1-dev: In the VAE category for Q2, three generators already exist in Synthbuster. We generate additional large-scale test sets from SD3.5 and Flux-1-dev, with 10,000 fake images each. Prompts are short descriptions generated by an LLM. Since corresponding real images are not available, we use the full PASCAL VOC training split, which contains 9,600 real images to compute balanced accuracy for the binary classes.
  • Satellite testing set (Q4): The real images comprise all images from four categories of the fMoW validation set (https://github.com/fMoW/dataset): amusement park (2,556 images) , car dealership (2,156 images), electric substation (2,626 images ), and stadium (2,575 images). Each real image has associated metadata including location (country). Fake images are generated using prompts of the form: “a fMoW satellite image of {class_name} in {country}”. The number of generated images matches the number of real images for each class.